# The Born rule – 100 years ago and today

### v2

@@ 2.3 Paradoxes and measurement context in Born’s rule
Already in 1926, Pauli [62, p.347] noticed the paradox that his objective position probability and Born’s objective momentum probability interpretation following from a spectral interpretation of the objective expectation form (BR-OE) of the Born rule are mutually incompatible. He expressed it in the following words: ”Man kann die Welt mit dem $p$-Auge und man kann sie mit dem $q$-Auge ansehen, aber wenn man beide Augen zugleich aufmachen will, dann wird man irre”
The 1927 paper by Jordan [37, p.811] cites Pauli [61] and extends Born’s second assumption further to an objective, measurement independent probability interpretation of inner products (probability amplitudes) of eigenstates of two arbitrary operators in place of position and momentum. without being aware of the conceptual problem this objective view poses when applied to noncommuting operators.
@@
Von Neumann’s change of perspective introduced the prominent reference to ”knowledge” (p.247). It would be interesting to have a thorough historical analysis of the meaning and use of the term knowledge in quantum mechanics. In these days, this term was not understood in a subjective way, but as the objective (through thought experiments theoretically accessible) knowledge of what is real and in principle observable about the system, whether observed or not.
This knowledge is encoded into a particular description, hence is an objective property of the description used to model the system. It is what the particular model claims to know, independent of who ”knows” or ”uses” it. Thus nothing mental is implied. A change of knowledge is therefore just a change of the details in the model used to describe a particular system. In particular, the situation regarding knowledge is precisely the same as in classical mechanics, where the models also depend on this kind of knowledge (or prejudice, if the model is chosen based on insufficient knowledge). Subjective knowledge seriously entered physics only much later, in 1957, with the ”subjective statistical mechanics” of Jaynes [34].
Later discussions of the statistical interpretation gradually switched from the objective versions of the Born rule to the measurement-based versions. As we have seen, de Broglie maintained in 1929 still both variants. This transition process was apparently completed in 1930 with the appearance of the very influential book by Dirac [21]. A review of the book, by Lennard-Jones [43], states: ”The book contains Dr. Dirac’s philosophy of the relation of theoretical and experimental physics. He believes that the main object of theory is to determine the possible results of an experiment, and to determine the probability that any one of these results will actually occur under given conditions. He regards it as quite unnecessary that any satisfying description of the whole course of the phenomena should be given. A mathematical machine is set up, and without asserting or believing that it is the same as Nature’s machine, we put in data at one end and take out the results at the other. As long as these results tally with those of Nature, (with the same data or initial conditions) we regard the machine as a satisfying theory. But so soon as a result is discovered not reproduced by the machine, we proceed to modify the machine until it produces the new result as well.
@@ 2.4 Collapse (state reduction)
Tied to the measurement formulations of the Born rule is a controversial collapse (or state reduction) postulate about the state after a measurement. State reduction was introduced informally in Heisenberg [31, p.186] (”Jede Ortsbestimmung reduziert also das Wellenpaket wieder auf seine ursprüngliehe Grösse”) and formally established by the authority of Dirac in several editions of his book:
”The state of the system after the observation must be an eigenstate of [the observable] $\alpha$, since the result of a measurement of $\alpha$ for this state must be a certainty.” (Dirac [21, p.49], first edition, 1930)
”Thus after the first measurement has been made, the system is in an eigenstate of the dynamical variable $\xi$, the eigenvalue it belongs to being equal to the result of the first measurement. This conclusion must still hold if the second measurement is not actually made. In this way we see that a measurement always causes the system to jump into an eigenstate of the dynamical variable that is being measured, the eigenvalue this eigenstate belongs to being equal to the result of the measurement.” (Dirac [22, p.36], third edition, 1936)
Dirac’s statement was discredited (without much success) in 1958 by another authority, Landau & Lifshitz [42, Section 7], who remark that the state after the measurement is in general not an eigenstate, This is apparent in the modern notion of a quantum process (see, e.g., Chuang & Nielsen [16] or Mohseni et al. [48]), which gives the correct post-measurement state for arbitrary POVM measurements.
@@
It would be interesting to have a more complete history of the collapse postulate and the extent of its validity.
@@ 3 The Born rule today
The most detailed, universal form of the Born rule is a very complex statement requiring considerable mathematical maturity to be fully understood. It is strange that such a complicated rule should be part of the basis of quantum theory.
@@
For this and other reasons, there were a host of attempts to derive the Born rule (in some measured form) from more natural assumptions; see Vaidman [68] for a recent survey. Schlosshauer & Fine [65] mention as desirable that a derivation of the Born rule gives ”physical insight into the emergence of quantum probabilities and the Born rule”.
@@
All attempts to derive the Born rule are doomed to failure if they do not define, in terms of the formal apparatus of quantum mechanics, the meaning of measurement, a concept appearing in the Born rule but not in the remaining basic rules of quantum mechanics. But measurement is usually introduced tacitly by invoking a seemingly self-evident consequence of the Born rule. Therefore the derivations found in the literature are usually circular and hence of questionable value.
@@
For example, Schonfeld [66] assumes in his elementary (but somewhat heuristic) ’derivation’ of Born’s rule on p.4 the seemingly self-evident fact that the thermal randomness of subcritical droplet formation ”exists independent of any measurement axioms, i.e., independent of whether anyone observes it.” While his description gives some insight into the origin of the Born rule, the problem of deriving how random nucleation (rather than a mixture of all possible mucleations) arises from a quantum statistical model (cf. Subsection 3.5) is not even touched. The traditional statistical account of nucleation by Langer [41] is purely classical. A quantum derivation of nucleation was attempted, e.g., by Lombardo et al. [44]. But (like many others) they assume without discussion that (under reasonable conditions) the ”Wigner function can then be interpreted as a classical probability distribution for coordinates and momenta”, an assertion that can be maintained as referring to a statistical distribution only by invoking the condensed version (BR-C) of the Born rule.
@@
Following Neumaier & Westra [54, Part II], we now give precise definitions for quantum expectations, statistical expectations, and quantum measurement devices, and derive from it the formulations (BR-POVM) and (BR-C) of the Born rule.
3.1 The detector response principle
@@
the intensity of the source. The intensity is nonnegative since $\rho$ is positive semidefinite. $\rho=0$ defines the empty state; it is the only state with zero intensity. Note the slight333 In contrast to our convention to normalize the trace of $\rho$ to the intensity, it is customary in the literature to normalize the intensity to be $1$. This can be achieved by dividing $\rho$ and all quantum values by the intensity. A disatvantage of this normalization is that $\rho$ loses one degree of freedom and the intensity of the source is no longer represented by the state. difference compared to conventional density operators, where the trace is instead fixed to be one.
@@
Since experimentally realizable detectors always produce only a finite number of possible results, we make the following definition. A quantum measurement device (in the following just called a measurement device) is characterized by a collection of detection elements labelled by labels $k$ from a finite set $K$ satisfying the detector response principle, given by the following postulate.
(DRP): A detection element $k$ responds to an incident stationary444 In order that a measured mean rate has a sensible operational meaning, the source must be reasonably stationary, at least during the time measurements are taken. source with density operator $\rho$ with a nonnegative mean rate $p_{k}$ depending linearly on $\rho$. The mean rates sum to the intensity of the source. Each $p_{k}$ is positive for at least one density operator $\rho$.
@@
As shown in Neumaier & Westra [54, Part II], the DRP leads naturally to all basic concepts and properties of modern quantum mechanics. In particular, it gives a precise operational meaning to quantum states, quantum detectors, quantum processes, and quantum instruments. This gives a perspective on the foundations of quantum mechanics that is quite different from the well-trodden path followed by most quantum mechanics textbooks.
@@ 3.2 The statistical interpretation of quantum mechanics
The key result for the theory of quantum measurements is the following detector response theorem, proved in [54, Section 4.1.2].
@@ 3.3 Derivation of the Born rule
If the intensity is normalized to one then $\sum p_{k}=1$, so that the normalized mean rate $p_{k}$ can be interpreted as the response probability of detector element $k$ given some response, or as the detection probability for the $k$th detection event. This gives an intuitive meaning for the $p_{k}$ in case of low intensity measurements. Formula (8), derived here from very simple first principles, becomes (4), but with $P_{k}$ not restricted to projections. Thus we have proved the POVM form (BR-POVM) of the Born rule. It gives the theoretical quantum values $\langle P_{k}\rangle$ a statistical interpretation as response probabilities $p_{k}$ of a quantum measurement device. Note that all probabilities are classical probabilities in the sense of Kolmogorov [38], as used everywhere in statistics, and can be approximated by relative frequencies with an accuracy given by the law of large numbers.
@@
It is shown in [54, Part II] that the DRP together with techniques from quantum tomography allow one to reconstruct from the DRP the complete edifice of quantum mechanics, including its dynamical and spectral consequences. In particular, the DRP is precisely what is needed to obtain the concepts and results from quantum information thoery.
@@
The orthogonality relations imply that $XP_{k}=x_{k}P_{k}$. Since the $P_{k}$ sum up to $1$, any $\psi\in\mathbb{H}$ can be decomposed into a sum $\psi=\displaystyle\sum P_{k}\psi=\sum\psi_{k}$ of vectors $\psi_{k}:=P_{k}\psi$ satisfying the equation $X\psi_{k}=XP_{k}\psi=x_{k}P_{k}\psi=x_{k}\psi_{k}$. Therefore each $\psi_{k}$ (if nonzero) is an eigenvector of $X$ (or of each component of $X$ in case the $x_{k}$ are not just numbers) corresponding to the eigenvalue $x_{k}$ of $X$. Since $P_{k}^{2}=P_{k}=P_{k}^{*}$, the $P_{k}$ are orthogonal projectors to the eigenspaces of the $x_{k}$.

### v3

@@ 2.3 Paradoxes and measurement context in Born’s rule
Already in 1926, Pauli [67, p.347] noticed the paradox that his objective position probability and Born’s objective momentum probability interpretation following from a spectral interpretation of the objective expectation form (BR-OE) of the Born rule are mutually incompatible. He expressed it in the following words: ”One can view the worls with the $p$-eye and one can view it with the $q$-eye, but if one wants to open both eyes at the same time, one gets crazy”; German original: ”Man kann die Welt mit dem $p$-Auge und man kann sie mit dem $q$-Auge ansehen, aber wenn man beide Augen zugleich aufmachen will, dann wird man irre”
The 1927 paper by Jordan [41, p.811] cites Pauli [65] and extends Born’s second assumption further to an objective, measurement independent probability interpretation of inner products (probability amplitudes) of eigenstates of two arbitrary operators in place of position and momentum. without being aware of the conceptual problem this objective view poses when applied to noncommuting operators.
@@
Later discussions of the statistical interpretation gradually switched from the objective versions of the Born rule to the measurement-based versions. As we have seen, de Broglie still maintained both variants in 1929. This transition process was apparently completed in 1930 with the appearance of the very influential book by Dirac [24]. A review of the book, by Lennard-Jones [47], states: ”The book contains Dr. Dirac’s philosophy of the relation of theoretical and experimental physics. He believes that the main object of theory is to determine the possible results of an experiment, and to determine the probability that any one of these results will actually occur under given conditions. He regards it as quite unnecessary that any satisfying description of the whole course of the phenomena should be given. A mathematical machine is set up, and without asserting or believing that it is the same as Nature’s machine, we put in data at one end and take out the results at the other. As long as these results tally with those of Nature, (with the same data or initial conditions) we regard the machine as a satisfying theory. But so soon as a result is discovered not reproduced by the machine, we proceed to modify the machine until it produces the new result as well.
@@
2.4 Knowledge
Von Neumann’s change of perspective introduced the prominent reference to ”knowledge” (von Neumann [60, p.247]). It would be interesting to have a thorough historical analysis of the meaning and use of the term knowledge in quantum mechanics. In these days, this term was not understood in a subjective way, but as the objective (through thought experiments theoretically accessible) knowledge of what is real and in principle observable about the system, whether observed or not.
However, the usage of the term ”knowledge” created a tension that never since disappeared, and was verbalized most prominently by Einstein throughout his life. In his obituary on Einstein, Puli [66] wrote in 1959: ” […] even when I could not agree with Einstein’s views. ’Physics is after all the description of reality’ he said to me, continuing, with a sarcastic glance in my direction ’or should I perhaps say physics is the description of what one merely imagines?’ This question clearly shows Einstein’s concern that the objective character of physics might be lost through a theory of the type of quantum mechanics, in that as a consequence of its wider conception of the objectivity of an explanation of nature the difference between physical reality and dream or hallucination might become blurred.
The objectivity of physics is however fully ensured in quantum mechanics in the following sense. Although in principle, according to the theory, it is in general only the statistics of series of experiments that is determined by laws, the observer is unable, even in the unpredictable single case, to influence the result of his observation – as for example the response of a counter at a particular instant of time. Further, personal qualities of the observer do not come into the theory in any way – the observation can be made by objective registering apparatus, the results of which are objectively available for anyone’s inspection.”
(German original: ”[…] auch wenn ich Einsteins Ansichten nicht zustimmen konnte. ’Physik ist doch die Beschreibung des Wirklichen’, sagte er zu mir und fuhr mit einem sarkastischen Blick auf mich fort: ’oder soll ich vielleicht sagen, Physik ist die Beschreibung dessen, was man sich bloss einbildet?’ Diese Frage zeigt deutlich Einsteins Besorgnis, dass durch eine Theorie vom Typus der Quantenmechanik der objektive Charakter der Physik verloren gehen könnte, indem durch deren weitere Fassung der Objektivität einer Naturerklärung der Unterschied der physikalischen Wirklichkeit von Traum oder Halluzination veschwommen werden könnte.
Die Objektivität der Physik ist in der Quantenmechanik jedoch im folgenden Sinn voll gewahrt. Obwohl nach der Theorie im Prinzip im allgemeinen nur die Statistik von Versuchsreihen gesetzmässig bestimmt ist, kann der Beobachter auch im nicht voraussagbaren Einzelfall das Resultat seiner Beobachtung – wie zum Beispiel das Ansprechen eines Zählers in einem bestimmten Zeitmoment – nicht beeinflussen.
Auch gehen persönliche Eigenschaften des Beobachters in keiner Weise in die Theorie ein, vielmehr kann die Beobachtung durch objektive Registrierapparate erfolgen, deren Resultate allen zur Einsicht objektiv vorliegen.”
Subjective knowledge (that Einstein hinted at with his question ”or should I perhaps say physics is the description of what one merely imagines?”) seriously entered physics only two years earlier, in 1957, with the ”subjective statistical mechanics” of Jaynes [38], thus aggravating the tension.
It is easy to see that the amount of knowledge needed to apply correctly the maximum entropy principle must be an intrinsic property of the system modeled. For example, take as system a turbulent fluid. If one knows only the total energy and the volume, and applies the maximum entropy principle, one gets a Helmholtz ensemble describing a fluid in equilibrium, not the turbulent fluid. Thus in spite of having used valid knowledge, the maximum entropy principle gives completely wrong results.
To correctly describe the latter in second quantization with the maximum entropy principle, it is necessary and sufficient that we apply it to an (approximate) knowledge of (at least) the quantum expectation values of the field operators for currents and densities of the fluid at all points in the support of the fluid.
But for a turbulent fluid one cannot get this (very detailed and very quickly changing) knowledge from observation, hence cannot know it in any meaningful sense! Instead, in all quantum derivations of fluid mechanical equations, one simply assumes that this knowledge (or rather the corresponding maximum entropy state) exists, and then draw consequences through statistical mechanics.
Thus the knowledge required to fix the state in nonequilibrium statistical mechanics is assumed knowledge, to be checked by its consequences. This makes sense only if the state (and hence any quantum expectation value) assumed in the model is an intrinsic property of the fluid, at least to the extent that the model predictions match observation.
Thus knowledge is encoded into a particular description, an objective property of the description used to model the system. It is what the particular model claims to know, independent of who ”knows” or ”uses” it. A change of knowledge is therefore just a change of the details in the model used to describe a particular system. In particular, the situation regarding knowledge is precisely the same as in classical mechanics, where the models also depend on this kind of knowledge (or prejudice, if the model is chosen based on insufficient knowledge).
Therefore, as Pauli had emphasized, nothing personal is implied in the term ”knowledge”. There is no more subjectivity in quantum statistical mechanics than in classical engineering physics: The only (and very restrictive) freedom a subject striving for maximal predictivity has is to choose a model correctly describing the system under study. Once the model is chosen, it determines the quality of the predictions, and, by comparison with corresponding observations, the quality of the knowledge built into the model. If the quality of this knowledge is insufficient, the model is objectively wrong, and was so even before the observations where made.
@@ 2.5 Collapse (state reduction)
Tied to the measurement formulations of the Born rule is a controversial collapse (or state reduction) postulate about the state after a measurement. State reduction was introduced informally in Heisenberg [35, p.186] (”Every determination of position therefore reduces the wave packet again to its original size”; German original: ”Jede Ortsbestimmung reduziert also das Wellenpaket wieder auf seine ursprüngliehe Grösse”) and formally established by the authority of Dirac in several editions of his book:
”The state of the system after the observation must be an eigenstate of [the observable] $\alpha$, since the result of a measurement of $\alpha$ for this state must be a certainty.” (Dirac [24, p.49], first edition, 1930)
”Thus after the first measurement has been made, the system is in an eigenstate of the dynamical variable $\xi$, the eigenvalue it belongs to being equal to the result of the first measurement. This conclusion must still hold if the second measurement is not actually made. In this way we see that a measurement always causes the system to jump into an eigenstate of the dynamical variable that is being measured, the eigenvalue this eigenstate belongs to being equal to the result of the measurement.” (Dirac [25, p.36], third edition, 1936)
Dirac’s statement was discredited (rightly, but without much success) in 1958 by another authority, Landau & Lifshitz [46, Section 7], who remark that the state after the measurement is in general not an eigenstate, This is apparent in the modern notion of a quantum process (see, e.g., Chuang & Nielsen [18] or Mohseni et al. [52]), which gives the correct post-measurement state for arbitrary POVM measurements.
@@
It would be interesting to have a more thorough history of the collapse postulate and the extent of its validity.
@@ 3 The Born rule today
The most detailed, universal form of the Born rule is a very complex statement requiring considerable mathematical maturity to be fully understood. Schlosshauer & Fine [70] mention as desirable that a derivation of the Born rule gives ”physical insight into the emergence of quantum probabilities and the Born rule”. It is indeed strange that such a complicated rule should be part of the basis of quantum theory.
@@
For this and other reasons, there were a host of attempts to derive the Born rule (in some measured form) from more natural assumptions; see Allahverdyan et al. [3, Chapter 2] for a thorough discussion of the state of the art until 2013, and Vaidman [73] (2020) for a more recent survey.
@@
All attempts to derive the Born rule are doomed to failure if they do not define, in terms of the formal apparatus of quantum mechanics, the meaning of measurement, a concept appearing in the Born rule but not in the remaining basic rules of quantum mechanics. But measurement is usually introduced tacitly by invoking a seemingly self-evident consequence of the Born rule. For example, use of the maximum entropy principle in a derivation implies circularity since by the discussion in Subsection 1.2, its application silently invokes (BR-C).
Another frequent mistake is to derive a probability distribution matching that of Born’s rule, without any actual measurement (with definite results) being involved. Only performed calculations within the theoretical machinery are performed, and only handwaving is used to claim that these calculations have anything to do with measurement. This was the case in early, decoherence-based ’derivations’ now known (Schlosshauer [69]) to lack an essential ingredient, the solution of the ”unique outcome problem”.
Therefore all derivations (known to me) found in the literature are either circular or unrelated to measurement, and hence of questionable value.
As examples, we consider three proposed derivations that appeared in the last few years.
@@
Schonfeld [71] (2021) assumes in his elementary (but somewhat heuristic) ’derivation’ of Born’s rule on p.4 the seemingly self-evident fact that the thermal randomness of subcritical droplet formation ”exists independent of any measurement axioms, i.e., independent of whether anyone observes it.” While his description gives some insight into the origin of the Born rule, the problem of deriving how random nucleation (rather than a mixture of all possible mucleations) arises from a quantum statistical model (cf. Subsection 3.5) is not even touched. The traditional statistical account of nucleation by Langer [45] is purely classical. A quantum derivation of nucleation was attempted, e.g., by Lombardo et al. [48]. But (like many others) they assume without discussion that (under reasonable conditions) the ”Wigner function can then be interpreted as a classical probability distribution for coordinates and momenta”, an assertion that can be maintained as referring to a statistical distribution only by invoking the condensed version (BR-C) of the Born rule.
Chanda & Bhattacharyya [17] (2021) claim on p.2 the emergence of Born’s rule in ”a dynamical model without an explicit invocation of an apparatus”. Thus they fall into the second trap menitioned above, that nothing ever is measured in their analysis. (In addition, they assume in their derivation not the unitary dynamics of a closed system but the dissipative Lindblad dynamics, which would have to be derived without using Born’s rule.)
Aharonov & Shushi [1] (2024) explicitly assume in their derivation – in the line after (1) – that measurement outcomes are eigenvalues, which is part of all spectral forms of the Born rule, and therefore should be proved in a complete derivation. In the line after (9), an observed value comes out of the blue by simple handwaving. (In addition, they explicitly assume a postulate, detailed after (12), which makes claims about unobservable properties of a particle between measurements – without even spelling out what a property should be. Thus instead of Born’s rule with its clear operational meaning they have to postulate an obscure assumptions unlikely to be acceptable to many physicists.)
3.1 The detector response principle
@@
Following Neumaier & Westra [58, Part II], we give precise definitions for quantum expectations, statistical expectations, and quantum measurement devices, and derive from these the formulations (BR-POVM) and (BR-C) of the Born rule.
@@
the intensity of the source. The intensity is nonnegative since $\rho$ is positive semidefinite. $\rho=0$ defines the empty state; it is the only state with zero intensity. Note the slight difference compared to conventional density operators, where the trace is instead fixed to be one. (In contrast to our convention to normalize the trace of $\rho$ to the intensity, it is customary in the literature to normalize the intensity to be $1$. This can be achieved by dividing $\rho$ and all quantum values by the intensity. A disatvantage of this normalization is that $\rho$ loses one degree of freedom and the intensity of the source is no longer represented by the state.)
@@
Since experimentally realizable detectors always produce only a finite number of possible results, we make the following definition. A quantum measurement device (in the following just called a measurement device) is characterized by a collection of detection elements labelled by labels $k$ from a finite set $K$ satisfying the detector response principle, given by the following postulate, (In order that a measured mean rate has a sensible operational meaning, the source must be reasonably stationary, at least during the time measurements are taken.)
(DRP): A detection element $k$ responds to an incident stationary source with density operator $\rho$ with a nonnegative mean rate $p_{k}$ depending linearly on $\rho$. The mean rates sum to the intensity of the source. Each $p_{k}$ is positive for at least one density operator $\rho$. (A slightly less precise version is already in my unpublished 2019 manuscript [54, p.8].)
@@
As shown in Neumaier & Westra [58, Part II], the DRP leads naturally to all basic concepts and properties of modern quantum mechanics. In particular, it gives a precise operational meaning to quantum states, quantum detectors, quantum processes, and quantum instruments. This gives a perspective on the foundations of quantum mechanics that is quite different from the well-trodden path followed by most quantum mechanics textbooks.
@@ 3.2 The statistical interpretation of quantum mechanics
The key result for the theory of quantum measurements is the following detector response theorem, proved in [58, Section 4.1.2].
@@ 3.3 Derivation of the Born rule
If the intensity is normalized to one then $\sum_{k}p_{k}=1$, so that the normalized mean rate $p_{k}$ can be interpreted as the response probability of detector element $k$ given some response, or as the detection probability for the $k$th detection event. This gives an intuitive meaning for the $p_{k}$ in case of low intensity measurements. Formula (8), derived here from very simple first principles, becomes (4), but with $P_{k}$ not restricted to projections. Thus we have proved the POVM form (BR-POVM) of the Born rule. It gives the theoretical quantum values $\langle P_{k}\rangle$ a statistical interpretation as response probabilities $p_{k}$ of a quantum measurement device. Note that all probabilities are classical probabilities in the sense of Kolmogorov [42], as used everywhere in statistics, and can be approximated by relative frequencies with an accuracy given by the law of large numbers.
@@
It is shown in [58, Part II] that the DRP together with techniques from quantum tomography allow one to reconstruct from the DRP the complete edifice of quantum mechanics, including its dynamical and spectral consequences. In particular, the DRP is precisely what is needed to obtain the concepts and results from quantum information thoery.
@@
The orthogonality relations imply that $XP_{k}=x_{k}P_{k}$. Since the $P_{k}$ sum up to $1$, any $\psi\in\mathbb{H}$ can be decomposed into a sum $\psi=\displaystyle\sum_{k}P_{k}\psi=\sum_{k}\psi_{k}$ of vectors $\psi_{k}:=P_{k}\psi$ satisfying the equation $X\psi_{k}=XP_{k}\psi=x_{k}P_{k}\psi=x_{k}\psi_{k}$. Therefore each $\psi_{k}$ (if nonzero) is an eigenvector of $X$ (or of each component of $X$ in case the $x_{k}$ are not just numbers) corresponding to the eigenvalue $x_{k}$ of $X$. Since $P_{k}^{2}=P_{k}=P_{k}^{*}$, the $P_{k}$ are orthogonal projectors to the eigenspaces of the $x_{k}$.